# Understanding and Embracing Imperfection in Physical Learning Networks

### v1

@@ Abstract
Performing machine learning (ML) with analog instead of digital signals offers advantages in speed and energy efficiency, but component and measurement imperfections can make nonlinear analog networks difficult to train. As a result, most schemes involve a precise digital model, either to train alone or in tandem with experiments. Here we take a different perspective: working in the analog domain, we characterize the consequences of the inherent imperfection of a physical learning system and, ultimately, overcome them. We train an analog network of self-adjusting resistors—a contrastive local learning network (CLLN)—for multiple tasks, and observe limit cycles and characteristic scaling behaviors absent in ‘perfect’ systems. We develop an analytical model that captures these phenomena by incorporating an uncontrolled but deterministic bias into the learning process. Our results suggest that imperfections limit precision and erase memory of previous tasks by continuously modifying the underlying representation of all learned tasks, akin to representational drift in the brain. Finally, we introduce and demonstrate a system-agnostic training method that greatly suppresses these effects. Our work points to a new, scalable approach in analog learning, one that eschews precise modeling and instead thrives in the mess of real systems.
@@ Introduction
Digital neural networks are expressive and malleable mathematical functions comprised of many simple and tunable nonlinear functions. Training these networks is typically accomplished using iterative gradient-based methods, which implicitly rely on extremely precise modeling of the network and adjustment of its tunable parameters. In digital networks, achieving such precision is straightforward because the model is the network, and simulated adjustments can be calculated and performed using any number of bits. However, when this precision is intentionally removed, for example by biasing the updates, training typically fails to achieve an acceptable accuracy [1]. This highlights a problem with applying these training methods to networks that exist in the physical world, where models, as well as actual components, measurements, and adjustments, will invariably be imperfect at some level.
@@
This simulation-reality gap is a major obstacle for many forms of neuromorphic computing [2] and physical learning [3], fields aimed at leveraging physical processes or entire physical systems to perform machine learning tasks. These approaches offer enormous potential advantages in energy-efficiency, speed, and other factors like physical robustness [4], but cannot assume pristine digital conditions. As a result, experiments modeled after digital ML systems typically use simulation and/or incorporate digital processing [5, 6, 7, 8, 9], limiting their ability to scale up. Alternatively, frameworks have been proposed for training physical systems that minimize energy or power [10, 11, 12, 13, 14, 15, 16]. Here, physical perturbation endows real signals with gradient information, avoiding differentiation of a model. However, imperfect measurements and adjustments of network parameters have prevented experimental demonstrations [4, 17, 18, 16] from reaching the accuracy and expressivity of their simulated counterparts [12, 15, 19], even when digital training is incorporated [14, 15].
@@
Biology has proven that learning complex tasks at scale is possible. For example, brains learn and function despite imperfect components, environmental variations, and even significant damage. Understanding how the brain overcomes these hurdles is extremely difficult, though obviously it does not succeed by using a companion digital simulation of itself to calculate its adjustments. To understand and ultimately overcome the impact of imperfection on the training of physical networks, it is fruitful to focus on simple systems that learn without digital aid.
@@
Here we observe, characterize, and ultimately suppress the effects of component imperfection and variability in a physically-constructed self-learning system. Specifically, we study a Contrastive Local Learning Network [20] (CLLN), an analog network in which each edge is a variable nonlinear resistor that autonomously self-adjusts according to a local rule. Such systems learn precise global functionality without digital or central computation. In training our CLLN alternatively between two different tasks, we observe non-equilibrium steady-state dynamics instead of convergence to a fixed point as seen in simulations of CLLNs. We introduce a model amenable to theoretical analysis and show that these dynamics appear when any bias is present in the adjustment of parameters. Finally, we use these insights to develop a modified training protocol, “overclamping”, which suppresses the effects of bias and drastically improves experimental results on classification tasks. Because this technique embraces the inherent imperfection of analog systems—requiring no digital modeling, additional measurements, or hardware improvements—it is not system-specific and can be applied broadly to improve training of physical systems.
@@
Figure 1: Training for Two Tasks A Image of the Contrastive Local Learning Network (CLLN) decorated with chosen input (yellow) and output (white) node locations. The system output is the difference $O=O_{+}-O_{-}$ between two chosen nodes. Inverted inputs $V_{i}^{\prime}=V_{+}-V_{i}$ of the variable inputs $V_{1}$ and $V_{2}$ are also imposed. Constant inputs have values $V_{+}=0.435$ V and $V_{-}=0.018$ V. Each edge of the network is a self-adjusting nonlinear resistor. B Training protocol. Two tasks $\alpha$ (blue) and $\beta$ (red) are alternatively trained for time $\tau/2$ until reaching a steady state. C Binary classification. We treat each class as a separate task. Background color is system output, labels $L_{\pm}=\pm$63 mV, and circles are training data. Steady state results are shown for $t=\tau/2$ and $t=\tau$ for $\tau=$ 0.2 s and 3 s. D Average error for each task (class) $E_{\alpha}$, $E_{\beta}$ over time in steady state. E Regression with two data points, separated into two tasks. $V_{2}$ is held at $V_{+}/2$ for this task. Labels are squares, output is black line, and panels are paired as in C. F Limit cycle error as in D. G Combined steady state error $\overline{E}$ vs $\tau$ for classification (circles) and regression (squares) experiments. Example times from C-F are indicated. H Same as G but for the cycle span $D$.
Experimental Methods
@@
We analyze the dynamics of a Contrastive Local Learning Network (CLLN), an experimental system described extensively in [20] and briefly here. The system is an analog electronic network of self-adjusting nonlinear resistors, where inputs are represented by enforced voltage boundary conditions on select nodes, and physics performs the ‘computation’ of outputs. The CLLN implements the Coupled Learning (CL) rule [12] that specifies the adjustment of each nonlinear resistor based only on local voltage drops. This rule, closely related to Equilibrium Propagation [11] and Contrastive Hebbian Learning [10], has been used successfully to train physical systems to perform nontrivial tasks such as nonlinear classification and regression [20, 21].
@@
The implementation of CL for analog electronic networks is as follows. First, a set of nodes are selected as inputs and output(s). A desired input-output response (task) is then prescribed by selecting training data points $[\vec{V}_{i},L_{i}]$ where $\vec{V}_{i}=V_{1,i},V_{2,i},...$ are the voltages imposed at the input nodes and $L_{i}$ is the desired output node voltage (label). In this work we consider only tasks with a single output voltage, but our framework and findings easily extend to tasks with multiple outputs [20]. The system is then exposed to two electronic states: the Free state where the inputs $\vec{V}_{i}$ are imposed and the network produces output $O_{i}^{F}$, and the Clamped state where the inputs and ‘nudged’ output $O_{i}^{C}$,
@@
where $S\approx 8\times 10^{-4}\ (\textrm{V}\Omega)^{-1}$, $V_{T}\approx 0.7$ V, and $\overline{V}$ is the average voltage of the two nodes in contact with the edge. This final term is the source of nonlinearity in our system [20].
@@
where $t_{h}$ is the amount of time the system is allowed to charge the capacitor (learning ‘on’) and $A_{0}\approx 2.5~{}(\text{V ms})^{-1}$, determined by choice of electronic components, see Appendix C. Our experiment uses twin networks to separately but simultaneously impose the Free and Clamped boundary conditions, and the two networks are updated in tandem so that the gate voltages on the twin edges are always the same.
@@
Here, we explore the effects of physical imperfection by training the system alternatively for two tasks $\alpha$ and $\beta$. We choose a period $\tau$, training the system for task $\alpha$ for time $\tau/2$, then task $\beta$ for time $\tau/2$, and then repeating these two steps over and over until the gate voltages reach a steady state, as shown schematically in Fig. 1B. Note that when a task involves multiple data points, they are cycled in order, each exposed to the network for time $t_{h}=100~{}\mu$s each until $\tau/2$ has elapsed.
Experimental Results
Unsurprisingly, we find that for sufficiently fast switching between the tasks (small periods $\tau$), the system finds a steady state where errors and outputs are relatively constant, as shown for $\tau=0.2$ s in Fig. 1C and D. However, as $\tau$ grows, a clear cyclic behavior emerges, with error for the task not being trained growing as the other shrinks. Note that the error for the ‘off’ task continues to evolve even when the ‘on’ task error finds a steady value, clearly seen in the $\tau=3s$ example.
We repeat this cyclic protocol for a pair of far simpler (regression) tasks, where each task is a single desired input-output relationship (data point). The same trends emerge, with minimal movement at small $\tau$ but tradeoffs between task errors at larger $\tau$. Here again, the longer the system trains for one task, the more its performance degrades for the other.
We find that in steady state, the combined error across both tasks and half-periods, $\overline{E}$ (equations in Appendix A), follows a similar pattern in both the classification and regression tasks, as shown in Fig. 1G. Namely, for short and long periods, the error is independent of $\tau$, However, in an intermediate regime, $\overline{E}$ grows monotonically with $\tau$.
By tracking the gate voltage values $\mathbf{G}=(G_{1},G_{2},...)$, we characterize the distance the system moves in one half period through the cycle span $D$,
@@
All of the behaviors noted in Fig. 1 were unanticipated. For a more precise network (e.g. a simulation of the network) learning these simple tasks would not display any steady-state cycles, but only zero-error fixed points ($\overline{E}(\tau)=0$ and $D(\tau)=0$). In the next section, we will show that these dynamics stem from imperfections in our learning system.
@@
We now examine a much smaller experimental system that displays similar phenomena, but which can be matched with analytic theory. Specifically, we interrogate a network of two self-adjusting edges and one fixed edge, as shown schematically in Fig. 2A. This network has three inputs, two that we hold constant $V_{+}=0.435$ V, $V_{-}=0.02$ V and one that we treat as our variable input as $V_{1}=V_{+}$ or $V_{-}$. We analyze two scenarios of switching between single-data point tasks (Fig. 2B and C), wherein task $\alpha$ is the same in both cases. For each single-data point task, there is a line of solutions in $G$ space; these are indicated as the blue and red dotted lines in Fig. 2F and G for the tasks in B and C, respectively. The intersection point of these solution lines, $\mathbf{G}^{*}$, corresponds to the joint solution. Following our previous (alternating) training protocol, a ‘perfect’ system will always reach and remain at this point, regardless of the period $\tau$.
@@
Figure 2: A Schematic of 2+1 edge experiment. B, C Pairs of single data point tasks, differing only by the desired output for task $\beta$. Solid, dashed gray lines connect data points in B,C. D Combined steady state error $\bar{E}$ as a function of $\tau$ for task pairs in B (circles) and C (squares). Vertical lines indicate $\tau$ values shown in F and G. Model predictions with no free parameters are shown as solid black lines. E Same as D but for Cycle Span $D$. F, G Visualization of limit cycles in $G$-space for task pairs in B and C, respectively. Colored points are experimental data for periods $\tau=$ 20 (lower left) and 800 ms (large bow-tie). Faded red and blue arrows indicate cycle direction. Colored dotted lines are individual task solution lines and their intersection is the joint solution $\mathbf{G^{*}}$. Gray arrows indicate measured bias $\mathbf{B}$. Thick black lines indicate the parameter maximum $G_{\text{max}}$. Note that edge $G_{1}$ has been modified to induce a different bias for task C, see Appendix E for details.
Modeling Imperfection
@@
where $\mathbf{G}^{*}$ is one of potentially many minima of $\mathcal{C}$ (a learning solution) and $\mathbf{M}$ is proportional to its Hessian evaluated at that point, with components $M_{ij}=\gamma\partial^{2}\mathcal{C}/\partial G_{i}\partial G_{j}\Big{|}_{\mathbf{G}^{*}}$. Here we discuss the global properties of Eq. (6) and refer the reader to the Appendix D for details on the derivation and solutions.
@@
Matching Experiment and Theory
To test our model we compare its predictions to experimental results. To begin, we first measure the experimental bias $\mathbf{B}$ using the drift along solution lines, see Appendix E, and the learning rate $\gamma=A_{0}/S$ from single-edge tests, see Appendix C. For the linearized dynamics we choose solution points that are close to the observed experimental trajectories. For compatible tasks, the solution points are taken as the common solution: $\mathbf{G}^{*}_{\alpha}=\mathbf{G}^{*}_{\beta}=\mathbf{G}^{*}$. These experimental measurements determine all of the parameters in the model, which displays excellent quantitative agreement with experiment in the combined error $\overline{E}$ and cycle span $D$. The agreement between theory and experiment is less strong for limit cycles, as shown by the black lines in panels 2D to G. Limit cycle predictions (Fig. 2F,G, black lines) are especially sensitive to deviations from the model assumptions, such as the assumption of linear behavior and state-independent bias. Nonetheless, the model displays the same behavior and trends as the experiment (colored cycles).
Figure 3: A inset A single data point ($\beta$, red) and a two-data point ($\alpha$, blue) task. A Combined steady state error $\overline{E}$ as a function of $\tau$. Model prediction with no free parameters is shown as a solid black line. Dashed black line indicates model prediction for zero bias (‘perfect’) evolution. Vertical lines highlight $\tau$ values shown in C. B Cycle span $D$ as a function of $\tau$ with model predictions as in A. C Visualization of four limit cycles in $G$-space for ($\tau=$ 25 ms, 100 ms, 400 ms, and 1600 ms). Gray arrow indicates measured bias $\mathbf{B}$. Faded red and blue arrows indicate cycle direction. The red dotted line indicates the solution space of task $\beta$, the blue dotted lines the solution spaces of each data point of task $\alpha$, and their intersection is the solution for the full $\alpha$ task. The X and Y axes for all panels have equal scale, and all Y axes have the same limits.

### v2

@@ Abstract
Performing machine learning with analog signals offers advantages in speed and energy efficiency, but sensitivity to component and measurement imperfections often foils training without a system-specific companion digital model. Here we take a different perspective, accepting and characterizing these inherent imperfections and ultimately overcoming them without digital models. We train an analog network of self-adjusting resistors—a contrastive local learning network—for multiple tasks, and observe limit cycles and scaling behaviors that limit precision, erase memory of previous tasks, and are absent in ‘perfect’ systems. We develop an analytical model capturing these phenomena as a consequence of an uncontrolled learning bias continuously modifying the underlying representation of learned tasks, reminiscent of representational drift in the brain. Finally, we introduce and demonstrate a system-agnostic training method that greatly suppresses these effects. Our work points to a new, scalable analog approach that eschews precise modeling and instead thrives in the mess of real systems.
@@ I Introduction
Digital neural networks are expressive and malleable mathematical functions comprised of many simple and tunable nonlinear functions. Training these networks is typically accomplished using iterative gradient-based methods, which implicitly rely on extremely precise modeling of the network and adjustment of its tunable parameters. In digital networks, achieving such precision is straightforward because the model is the network, and simulated adjustments can be calculated and performed using any number of bits. However, when this precision is intentionally removed, for example by biasing the updates, training typically fails to achieve an acceptable accuracy [1]. This highlights a problem with applying these training methods to physically-instantiated analog networks, where modeling, measurement, and operation of components will invariably be imperfect at some level.
@@
This simulation-reality gap is a major obstacle for many forms of neuromorphic computing [2, 3] and physical learning [4], fields aimed at leveraging physical processes or entire physical systems to perform machine learning tasks. These approaches offer enormous potential advantages in energy-efficiency, speed, and other factors like physical robustness [5], but cannot assume pristine digital conditions. As a result, experiments modeled after digital ML systems typically use simulation and/or incorporate digital processing [6, 7, 8, 9, 10], limiting their ability to scale up. Alternatively, frameworks have been proposed for training physical systems that minimize energy or power [11, 12, 13, 14, 15, 16, 17]. Here, physical perturbation endows real signals with gradient information, avoiding differentiation of a model. However, imperfect measurements and adjustments of network parameters have prevented experimental demonstrations [5, 18, 19, 17] from reaching the accuracy and expressivity of their simulated counterparts [13, 16, 20], even when digital training is incorporated [15, 16].
@@
Biology has proven that learning complex tasks at scale is possible. For example, brains learn and function despite imperfect components, environmental variations, and even substantial damage. Understanding how the brain overcomes these hurdles is extremely difficult, though obviously it does not succeed by using a companion digital simulation of itself to calculate its adjustments. To understand and ultimately overcome the impact of imperfection on the training of physical networks, it is fruitful to focus on simple systems that learn without digital aid.
@@
Here we observe, characterize, and ultimately suppress the effects of component imperfection and variability in a physically-constructed self-learning system. Specifically, we study a Contrastive Local Learning Network [21] (CLLN), an analog network in which each edge is a variable nonlinear resistor that autonomously self-adjusts according to a local rule. Such systems learn precise global functionality without digital or central computation. In training our CLLN alternatively between two different tasks, we observe non-equilibrium steady-state dynamics instead of convergence to a fixed point as seen in simulations of CLLNs. We introduce a model amenable to theoretical analysis and show that these dynamics appear when any bias is present in the adjustment of parameters. Importantly, biases in measurement, calculation, and application are generic and unavoidable features of analog physical hardware [3]. Finally, we use these insights to develop a modified training protocol, “overclamping”, which suppresses the effects of bias (and noise, though negligible in our experiments) and drastically improves experimental results on classification tasks. Because this technique embraces the inherent imperfection of analog systems—requiring no digital modeling, additional measurements, or hardware improvements—it is not system-specific and can be applied broadly to improve training of physical systems.
@@
Figure 1: Training for Two Tasks A Image of the Contrastive Local Learning Network (CLLN) decorated with chosen input (yellow) and output (white) node locations. The network is a grid of self-adjusting nonlinear resistors – specifically n-channel enhancement MOSFET transistors – and nodes are simply connections between edges. Two copies of the network are constructed (they coexist on the same breadboards) to run the Free and Clamped states (see text). The system output is the difference $O=O_{+}-O_{-}$ between two chosen nodes, and variable inputs $V_{1}$ and $V_{2}$, their inverses $V_{i}^{\prime}=V_{+}-V_{i}$, and two constant values $V_{+}=0.435$ V and $V_{-}=0.018$ V are imposed as voltage boundary conditions. The differential output and inverted inputs both give the system access to negative relationships between input and output, permitting more complex functionality. B Training protocol. Two tasks $\alpha$ (blue) and $\beta$ (red) are alternatively trained for time $\tau/2$ until reaching a steady state. C Binary classification. We treat each class as a separate task. Background color is system output, labels $L_{\pm}=\pm$63 mV, and circles are training data. Steady state results are shown for $t=\tau/2$ and $t=\tau$ for $\tau=$ 0.2 s and 3 s. D Average error for each task (class) $E_{\alpha}$, $E_{\beta}$ over time in steady state. E Regression with two data points, separated into two tasks. $V_{2}$ is held at $V_{+}/2$ for this task. Labels are squares, output is black line, and panels are paired as in C. F Limit cycle error as in D. G Combined steady state error $\overline{E}$ vs $\tau$ for classification (circles) and regression (squares) experiments. Example times from C-F are indicated. H Same as G but for the cycle span $D=\lVert\mathbf{G}(\tau/2)-\mathbf{G}(\tau)\rVert$, that is, the straight-line distance the system travels in a half-cycle.
II Results
II.1 Description of the experimental setup
@@
We analyze the dynamics of a Contrastive Local Learning Network (CLLN), an experimental system described extensively in [21] and briefly here. The system is an analog electronic network of self-adjusting nonlinear resistors, where inputs are represented by enforced voltage boundary conditions on select nodes, and physics performs the ‘computation’ of outputs by enforcing Kirchhoff’s current and voltage laws. The CLLN implements the Coupled Learning (CL) rule [13] that specifies the adjustment of each nonlinear resistor based only on local voltage drops. This rule, closely related to Equilibrium Propagation [12] and Contrastive Hebbian Learning [11], has been used successfully to train physical systems to perform nontrivial tasks such as nonlinear classification and regression [21, 22].
@@
The implementation of CL for analog electronic networks is as follows. First, a set of nodes are selected as inputs and output(s). A desired input-output response (task) is then prescribed by selecting training data points $\vec{V}_{i},L_{i}]$ where $\vec{V}_{i}=V_{1,i},V_{2,i},$... are the voltages imposed at the input nodes and $L_{i}$ is the desired output node voltage (label). In this work we consider only tasks with a single output voltage, but our framework and findings easily extend to tasks with multiple outputs [21]. The system is then exposed to two electronic states: the Free state where the inputs $\vec{V}_{i}$ are imposed and the network produces output $O_{i}^{F}$, and the Clamped state where the inputs and ‘nudged’ output $O_{i}^{C}$,
@@
where $S\approx 8\times 10^{-4}\ (\textrm{V}\text{ ohms})^{-1}$, $V_{T}\approx 0.7$ V, and $\overline{V}$ is the average voltage of the two nodes in contact with the edge. Despite being in the Ohmic regime, Eq. (2) leads to a non-linear relation between the current across the transistor and its terminal voltages, as the conductance itself depends on the average voltage $\overline{V}$—this is the source of nonlinearity in our system [21]. We also emphasize that while our experimental network has a grid structure, it can accommodate any architecture. Moreover, unlike crossbar-based analog implementations [3], our CLLN does not rely on accurately implementing matrix-vector-multiplication needed for feed-forward networks; it is a fully recurrent network implementing local learning rules.
@@
where $t_{h}$ is the amount of time the system is allowed to charge the capacitor (learning ‘on’) and $A_{0}\approx 2.5\times 10^{3}~(\text{V s})^{-1}$, determined by choice of electronic components, see Materials and Methods. Our experiment uses twin networks to separately but simultaneously impose the Free and Clamped boundary conditions, and the two networks are updated in tandem so that the gate voltages on the twin edges are always the same.
@@
[…]
II.2 Experimental Results
Fig. 1C and D illustrate the alternate training in binary classification, where each class corresponds to a task. For classification, the network is trained using the normal protocol, but clamping is removed ($\eta\rightarrow 0$) when a datapoint is ‘very correctly classified’, here meaning that its output is beyond the label on the correct side of the boundary. This effectively creates a shifted hinge loss, see Materials and Methods for details. Unsurprisingly, we find that for sufficiently fast switching between the tasks (small periods $\tau$), the system finds a steady state where errors (squared hinge loss) and outputs are relatively constant, as shown for $\tau=0.2$ s. However, as $\tau$ grows, a clear cyclic behavior emerges, with error for the task not being trained growing as the other shrinks. Note that the error for the ‘off’ task continues to evolve even when the ‘on’ task error finds a steady value, clearly seen in the $\tau=3s$ example.
We repeat this cyclic protocol for a pair of far simpler (regression) tasks, where each task is a single desired input-output relationship (data point). The same trends emerge, with minimal movement at small $\tau$ but tradeoffs between task errors (mean-squared errors) at larger $\tau$. Here again, the longer the system trains for one task, the more its performance degrades for the other, resembling the phenomenon of catastrophic forgetting in machine learning [23]—where memories of old tasks are erased by learning new ones—although here it is driven by physical imperfections, rather than training schedule alone.
We find that in steady state, the combined error across both tasks and half-periods, $\overline{E}$ (equations in Materials and Methods), follows a similar pattern in both the classification and regression tasks, as shown in Fig. 1G. Specifically, the error is independent of the training period $\tau$ for both short and long cycles, but in an intermediate regime it increases monotonically with $\tau$.
[…]
@@ II.2 Experimental Results
All of the behaviors noted in Fig. 1 were unanticipated. For a more precise network (e.g. a simulation of the network) learning these simple tasks would not display any steady-state cycles, but only zero-error fixed points ($\overline{E}(\tau)=0$ and $D(\tau)=0$). Notably, the transition out of the low-error plateau occurs as $\tau$ eclipses the characteristic timescale of learning ($\tau_{0}\approx 18$ ms, set by chosen resistors and capacitors [21]) times the number of datapoints. This suggests the system ‘sees’ datapoints collectively at small $\tau$ and individually at large $\tau$. However, in a simulation, this would still only affect the transient behavior prior to finding the zero-error solution. In the next section, we will show that these dynamics stem from imperfections in our learning system.
@@
We now examine a much smaller experimental system that displays similar phenomena, but which can be matched with analytic theory. Specifically, we interrogate a network of two self-adjusting edges and one fixed edge, as shown schematically in Fig. 2A. This network has three inputs, two that we hold constant $_{+}=0.435 V, V$$_{-}=0.02 V$ and one that we treat as our variable input as $V_{1}=V_{+}$ or $V_{-}$. We analyze two scenarios of switching between single-data point tasks (Fig. 2B and C), wherein task $\alpha$ is the same in both cases. For each single-data point task, there is a line of solutions in $G$ space; these are indicated as the blue and red dotted lines in Fig. 2F and G for the tasks in B and C, respectively. The intersection point of these solution lines, $\mathbf{G}^{*}$, corresponds to the joint solution. Following our previous (alternating) training protocol, a ‘perfect’ system will always reach and remain at this point, regardless of the period $\tau$.
@@
Figure 2: Learning Dynamics in a Small Experimental Network A Schematic of 2+1 edge experiment. B, C Pairs of single data point tasks, differing only by the desired output for task $\beta$. Solid, dashed gray lines connect data points in B,C. D Combined steady state error $\bar{E}$ as a function of $\tau$ for task pairs in B (circles) and C (squares). Vertical lines indicate $\tau$ values shown in F and G. Model predictions with no free parameters are shown as solid black lines. E Same as D but for Cycle Span $D$. F, G Visualization of limit cycles in $G$-space for task pairs in B and C, respectively. Colored points are experimental data for periods $\tau=$ 20 (lower left) and 800 ms (large bow-tie). Faded red and blue arrows indicate cycle direction. Colored dotted lines are individual task solution lines and their intersection is the joint solution $\mathbf{G^{*}}$. Gray arrows indicate measured bias $\mathbf{B}$. Thick black lines indicate the parameter maximum $G_{\text{max}}$. Note that edge $G_{1}$ has been modified to induce a different bias for task C, see Materials and Methods for details.
II.3 Modeling Imperfection
@@
where $\mathbf{G}^{*}$ is one of potentially many minima of $\mathcal{C}$ (a learning solution) and $\mathbf{M}$ is proportional to its Hessian evaluated at that point, with components $M_{ij}=\gamma\partial^{2}\mathcal{C}/\partial G_{i}\partial G_{j}\Big|_{\mathbf{G}^{*}}$. Here we discuss the global properties of Eq. (6) and refer the reader to the Materials and Methods for details on the derivation and solutions.
@@
II.4 Matching Experiment and Theory
To test our model we compare its predictions to experimental results. To begin, we first measure the experimental bias $\mathbf{B}$ using the drift along solution lines and the learning rate $\gamma=A_{0}/S$ from single-edge tests, see Materials and Methods. For the linearized dynamics we choose solution points that are close to the observed experimental trajectories. For compatible tasks, the solution points are taken as the common solution: $\mathbf{G}^{*}_{\alpha}=\mathbf{G}^{*}_{\beta}=\mathbf{G}^{*}$. These experimental measurements determine all of the parameters in the model, which displays excellent quantitative agreement with experiment in the combined error $\overline{E}$ and cycle span $D$. The agreement between theory and experiment is less strong for limit cycles, as shown by the black lines in panels 2D to G. Limit cycle predictions (Fig. 2F,G, black lines) are especially sensitive to deviations from the model assumptions, such as the assumption of linear behavior and state-independent bias. Nonetheless, the model displays the same behavior and trends as the experiment (colored cycles).
Figure 3: Incompatible Tasks a Small Experimental Network A inset A single data point ($\beta$, red) and a two-data point ($\alpha$, blue) task. A Combined steady state error $\overline{E}$ as a function of $\tau$. Model prediction with no free parameters is shown as a solid black line. Dashed black line indicates model prediction for zero bias (‘perfect’) evolution. Vertical lines highlight $\tau$ values shown in C. B Cycle span $D$ as a function of $\tau$ with model predictions as in A. C Visualization of four limit cycles in $G$-space for ($\tau=$ 25 ms, 100 ms, 400 ms, and 1600 ms). Gray arrow indicates measured bias $\mathbf{B}$. Faded red and blue arrows indicate cycle direction. The red dotted line indicates the solution space of task $\beta$, the blue dotted lines the solution spaces of each data point of task $\alpha$, and their intersection is the solution for the full $\alpha$ task. The X and Y axes for all panels have equal scale, and all Y axes have the same limits.